OpenStudy (waheguru):

how do i solve this 4= 2 (p-2) there is a negative sign in front of it - - 3

OpenStudy (waheguru):

NEATIVE ISGN IN FRON OF FRACTION

OpenStudy (anonymous):



OpenStudy (waheguru):

how u get -6 pmn??????

OpenStudy (anonymous):

\[4=\frac{-2}3(p-2)\] \[12=-2(p-2)\] \[\frac{-12}2=p-2\] \[-6=p-2\] \[-6+2=p\] \[-4=p\]

OpenStudy (anonymous):

i multiplyed by 3 then devided by -2.

OpenStudy (waheguru):

yea but when u get to -12= 2(p-2) dont you multiply the 2 so u get 2p-4??

OpenStudy (anonymous):

yes u can do like that also \[-12=2p-4\] \[-12+4=2p\] \[-8=2p\] \[-8/2=p\] \[-4=p\] answer is same

OpenStudy (waheguru):

plzz help me i am soo confused

OpenStudy (anonymous):

k whats confusing u ?

OpenStudy (waheguru):

this is how i did it plzz correct any errors 4= -2/3(three is also negative as the whole fraction is negative) (p-2) 4(-3)=-2(p-2) -12=2p+4 -12-4=-2p -16=-2p -16/-2=p 8=p ?????

OpenStudy (waheguru):

why a i getting differnt answers

OpenStudy (anonymous):

-2/3 is there there is no -2/-3 waheguru u did 4(-3) =-2(p-2) it should have been 4(3) = -2(p-2)

OpenStudy (anonymous):

got it now ? waheguru

OpenStudy (waheguru):

oh so the bottom number is never a negative?

OpenStudy (anonymous):

in a standard form it is not but a fraction 2/(-3) = -2/3 ok!!!

OpenStudy (waheguru):

ok thxxxxxxxxxxxxxxxxxxxx